OpenStudy (anonymous):

How to find equation of line from point (3, -5) with a slope of 1/5

OpenStudy (anonymous):

y=1/5x + b plug in (3,-5) to x and y and solve for b -5=1/5(3) +b b=-5 - 1/5(3)